# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 4. Preliminaries on Random Walks
$0<\log M,h\left(\varphi_{l}\right)<\mathfrak{X}(7)3^{59}((\log C_{1}+\log(2|S|))l+3\log 3)<C_{2}l$ (4.14)
where $C_{2}=3^{61}\mathfrak{X}(7)\log 2C_{1}|S|$.
Now we take $(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\in\mathbb{Z}^{7}$ such that
$(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\equiv(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}})(\operatorname{mod}Q^{\prime}).$
It follows that
$M\equiv\sum_{l=1}^{N}F_{l}\varphi_{l}\Bigg{|}_{(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})}\equiv 0\ (\operatorname{mod}{Q^{\prime}}),$
@@
$x_{1}+\tilde{Y}y_{1}+\tilde{Z}z_{1}+\tilde{W}w_{1}+\tilde{A}x_{2}+\tilde{B}y_{2}+\tilde{C}z_{2}+\tilde{D}w_{2}=0$
Rescaling to get rid of the common denominator, we obtain $(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Z}^{8}$ such that $\text{gcd}(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=1$, and for all $\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right)\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)\in\mathcal{G}$, we have
$\tilde{X}_{1}x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+\tilde{Y}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}=0.$
We finish the proof of Lemma 4.5 of the case $n=0$ by applying Lemma 4.7 to the linear form $\tilde{L}$ determined by the constants $\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}.$ For $1\ll_{S}l<c_{S}\log Q$,
$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})\mid L(g)\equiv 0(\operatorname{mod}{Q})\right\}\right)$

### v4

@@ 4. Preliminaries on Random Walks
$0<\log M,h\left(\varphi_{\gamma}\right)<\mathfrak{X}(8)3^{67}((\log N_{1}+\log(2|S|))l+\log Q_{1}+3\log 3)<N_{3}(l+\log Q_{1})$ (4.18)
where
$N_{3}=3^{69}\mathfrak{X}(8)\log(2|S|N_{1})$ (4.19)
Let $(z_{0},\cdots,z_{7})\in\mathbb{Z}^{8}$ such that $(z_{0},\cdots,z_{7})\equiv(\gamma_{0}\overline{n_{1}},\cdots,\gamma_{7}\overline{n_{1}})(\operatorname{mod}Q_{2})$, where $\overline{n_{1}}$ is the multiplicative inverse of $n_{1}(\operatorname{mod}Q_{2})$, then (4.17) implies
$\sum_{\gamma\in\Gamma}f_{\gamma}(z_{0},\cdots,z_{8})\varphi_{\gamma}(z_{0},\cdots,z_{8})\equiv M(\operatorname{mod}Q_{2}),$
so that
$M\geq Q_{2},$ (4.20)
which contradicts (4.18) if we take $c_{S}\leq\frac{1}{3N_{3}}\log Q$. Hence $\{f_{\gamma},\gamma\in\mathcal{G}\}$ have a common zero $(v_{1},\cdots,v_{8})\in\mathbb{Z}^{8}-\{\vec{\bf 0}\}$. Applying Lemma 4.7 then yields
$\chi_{S}^{(l)}(\mathcal{G}^{\prime})<e^{-cl},$ (4.21)
for some $c>0$.
Case 2.2: $Q_{1}\geq Q_{2}^{\frac{1}{N_{3}}}$. In this case the equation $L(\gamma)-n\equiv 0(\operatorname{mod}Q)$ implies $L(\gamma)\equiv 0(\operatorname{mod}Q_{1})$. Thus we reduce this case to Case 1. ∎
5. Bounded generation for simple factors
Let $\mathbb{P}_{0}$ be the projection from $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ to $\text{SL}_{2}(\mathbb{Z})$, and let $\mathbb{P}_{1}$ ($\mathbb{P}_{2}$, resp.) be the projection map from $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ to its first factor (second factor, resp.). The goal of this section is to prove the following bounded generation result for simple factors of $\Lambda(\operatorname{mod}q)$, where $q$ is a product of large powers of primes:
Proposition 5.1.
Let $S$ be a finite symmetric set in $\Lambda$ which generates a Zariski dense subgroup. There exist positive functions $\rho=\rho(\delta)=O_{S}(\delta),C=C(\delta)\in\mathbb{Z}_{+}$ defined for any sufficiently small $\delta>0$, such that the following holds: For any sufficiently large $q=\prod_{i\in I}p_{i}^{n_{i}},n_{i}\geq\frac{3}{\delta}$, and for any $l\in\mathbb{Z}_{+}$ and $A\subset\Lambda$ that satisfy
$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q$
and
$|\pi_{q}(A\cdot A\cdot A)|\leq|\pi_{q}(A)|^{1+\delta},$ (5.2)
we have
$\Gamma(q_{*}^{\{\rho\}})/\Gamma(q_{*})\subset\begin{cases}&(\mathbb{P}_{i}A)^{C}(\operatorname{mod}q_{*}),i=1,2,\hskip 8.53581pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}),\\ &(\mathbb{P}_{0}A)^{C}(\operatorname{mod}q_{*}),\hskip 45.5244pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}.\end{cases}$ (5.3)
where $q_{*}\|q$ and $q_{*}\geq q^{c}$ for some $c>0$ depending only on $S$.
We focus on the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and $i=1$. The proofs for the remaining two cases are virtually identical. Our proof follows the framework in [BG09], which are outlined below. These statements are understood to hold under the reduction of appropriate moduli.
  1. (a)
    Use non-concentration of random walks (Proposition 4.1) to find a large set of matrices with distinct traces from $\mathbb{P}_{1}(A^{2})$ (Lemma 5.7).
  2. (b)
    Assuming (5.2) holds, use the matrices obtained from (a) to create a large set of commutative matrices (Lemma 5.18).
  3. (c)
    Diagonalize the matrices from (b) over a quadratic field to generate a large set of diagonal matrices that satisfy multiple density and proximity-to-identity constraints (Lemma 5.32).
  4. (d)
    Apply the sum-product theorem (Theorem B) to create a line in $\mathfrak{sl}_{2}$ (Lemma 5.50).
  5. (e)
    Use random walks (Proposition 4.2) to move the line from (d) to different directions by conjugation, thus to create a thick segment in $\mathfrak{sl}_{2}$ (Lemma 5.92).
  6. (f)
    Iteratively take commutators of the segment obtained from (e) to finish the proof of Proposition 5.1.
Remark 5.4.
The method for finding commutative matrices can be traced back to [Hel08].
The presentation in this section involves several constants. For readers’ convenience we record them here.
Let $c_{1},c_{2}$ be the implied constants for $c$ in Propositions 4.1, 4.2. We take $c_{1},c_{2}<1$.
Let $C_{1},C_{2},C_{3},\varepsilon=\varepsilon_{0}$ be the implied constants in Theorem B for $\alpha=\frac{c_{1}}{6}$.
Let $\rho_{1},\rho_{2}$ be two small quantities linearly depending on the (only) variable $\delta$:
$\rho_{1}=L_{1}\delta,\hskip 28.45274pt\rho_{2}=L_{2}\rho_{1}=L_{1}L_{2}\delta.$ (5.5)
Here $L_{1}$ and $L_{2}$ are two big constants up to several adjustments in the proof process, and we assume at each adjustment, they satisfy not only the current requirement but also all preceding ones. While it is important $L_{1}$ and $L_{2}$ only depend on $S$, we take $L_{1}$ and $L_{2}$ so large that any positive linear combination of $\delta$ and $\rho_{1}$ appearing in our presentation is dominated by the $\rho_{1}$ term, and any positive linear combination of $\rho_{1}$ and $\rho_{2}$ is dominated by the $\rho_{2}$ term. So $\delta$ should be considered significantly smaller than $\rho_{1}$, which in turn is significantly smaller than $\rho_{2}$. The main parameter at (5.5) is $\rho_{2}$; the parameter $\rho_{1}$ plays a correction role.
Unless otherwise specified, the Landau notations $\Theta,\Omega,O$ in this section describe functions of $\delta$ and the implied constants depend only on $S$.
We also need to introduce two small quantities $w_{1},w_{2}=\Theta(\rho_{2})$ given in Lemma 5.32. So by our convention above, $w_{1}$, $w_{2}$ is much larger than $\rho_{1}$.
5.1. A large set of traces
Let
$A_{0}=(\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A))\cap\Gamma(\tilde{q})\subset\mathbb{P}_{1}(A^{2})\subset\text{SL}_{2}(\mathbb{Z}),$
where $\tilde{q}=\prod_{p|q}p$ is the square-free part of $q$.
Let $\nu$ be the push forward of $\chi_{S}$ under the projection map $\mathbb{P}_{1}$.
For each $x\in\Gamma_{\tilde{q}}$, let
@@ 5.1. A large set of traces
Then,
$\nu^{(2l)}(A_{0})$ $=\sum_{x\in\Gamma_{\tilde{q}}}\nu^{(l)}(B_{x})\nu^{(l)}(B_{x^{-1}})\stackrel{{\scriptstyle\text{symmetry of $A$}}}{{=}}\sum_{x\in\Gamma_{\tilde{q}}}\nu^{(l)}(B_{x})^{2}$
$\stackrel{{\scriptstyle\text{Cauchy-Schwarz}}}{{\geq}}\frac{\nu^{(l)}(\mathbb{P}_{1}(A))^{2}}{|\Gamma_{\tilde{q}}|}>q^{-2\delta}(\tilde{q})^{-3}\geq q^{-3\delta},$ (5.6)
where in the last inequality we have used that the exponent for each prime divisor of $q$ exceeds $\frac{3}{\delta}$.
The goal of this section is to prove
Lemma 5.7.
There is a set $A_{1}\subset A_{0}$, an exact divisor $q_{1}\|q$ with $q_{1}>q^{\frac{1}{2}}$, such that
$|A_{1}(\operatorname{mod}q_{1}^{\{\rho_{2}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{2}},$ (5.8)
$\forall\gamma_{1},\gamma_{2}\in A_{1},\gamma_{1}\neq\gamma_{2}\Rightarrow\text{tr}(\gamma_{1})\not\equiv\text{tr}(\gamma_{2})(\operatorname{mod}q_{1}^{\{\rho_{2}\}})$ (5.9)
$\forall\gamma=\left(\begin{matrix}\gamma_{11}&\gamma_{12}\\ \gamma_{21}&\gamma_{22}\end{matrix}\right)\in A_{1},\text{tr}(\gamma)\not\equiv\pm 2,\gamma_{21}\not\equiv 0(\operatorname{mod}p^{[4n\rho_{1}]}),\forall p^{n}\|q_{1},$ (5.10)
where $c_{1}$ is the implied constant $c$ from Proposition 4.1.
Remark 5.11.
Lemma 5.7 gives us a large set of matrices with distinct traces not close to $\pm 2$ at all non-Archimedean places $p$ where $v_{p}(q_{1})>0$ . We also need conditions (5.9) and (5.10) because they ensure that we can diagonalize any element from $A_{1}$ by a matrix with controlled denominator.
Proof of Lemma 5.7.
For each $Q|q$, let
$\mathcal{E}_{1}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\text{tr}(\gamma)\equiv\pm 2\equiv 0(\operatorname{mod}Q))\},$
Proposition 4.1 gives
$\nu^{(2l)}(\mathcal{E}_{1}(Q)),\nu^{(2l)}(\mathcal{E}_{2}(Q))<Q^{-c_{1}},$
for sufficiently large $Q$. Let
$A_{1}^{\prime}=A_{0}-\bigcup_{\begin{subarray}{c}Q|q\\ Q\geq q^{\rho_{1}}\end{subarray}}(\mathcal{E}_{1}(Q)\cup\mathcal{E}_{2}(Q)).$
Since the number of divisors of $q$ is bounded by $q^{0+}$, it follows that
$\nu^{(2l)}(A_{1}^{\prime})>q^{-3\delta-},$
if we let
$L_{1}>\frac{1}{c_{1}}.$ (5.12)
For each $\gamma\in A_{1}^{\prime}$, let
$q_{\gamma}^{{}^{\prime}}=\prod_{\begin{subarray}{c}p^{n}\|q\\ \text{tr}(\gamma)\not\equiv\pm 2(\operatorname{mod}p^{[4n\rho_{1}]})\end{subarray}}p^{n},$
and
$q_{\gamma}^{{}^{\prime\prime}}=\prod_{\begin{subarray}{c}p^{n}\|q\\ \gamma_{21}\not\equiv 0(\operatorname{mod}p^{[4n\rho_{1}]})\end{subarray}}p^{n}.$
Since $\text{gcd}(\text{tr}(\gamma)\pm 2,q)<q^{\rho_{1}}$ and $\text{gcd}(\gamma_{21},q)<q^{\rho_{1}}$, we must have $q_{\gamma}^{\prime},q_{\gamma}^{\prime\prime}>q^{\frac{3}{4}}$.
Let $q_{\gamma}=\text{gcd}(q_{\gamma}^{\prime},q_{\gamma}^{\prime\prime})$. Then $q_{\gamma}>q^{\frac{1}{2}}$, and for any $p^{n}\|q_{\gamma}$,
$\text{tr}(\gamma)\not\equiv\pm 2(\operatorname{mod}p^{[4\rho_{1}n]}),$ (5.13)
$\gamma_{21}\not\equiv 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.14)
Since again the number of divisors of $q$ (which bounds the number of choices for $q_{\gamma}$) is bounded by $q^{0+}$, there exists $q_{1}\|q,q_{1}>q^{\frac{1}{2}}$ and a set $A_{1}^{\prime\prime}\subset A_{1}^{\prime}$, with
$\nu^{(2l)}(A_{1}^{\prime\prime})>q^{-3\delta-}>q_{1}^{-6\delta-},$ (5.15)
such that for any $\gamma\in A_{1}^{\prime\prime}$, any $p^{n}\|q_{1}$, (5.13) and (5.14) are satisfied.
Let
$W(n):=\{\gamma\in\Gamma:\text{tr}(\gamma)=n(\operatorname{mod}q_{1}^{\{\rho_{2}\}})\}.$
By Proposition 4.1, for any $n\in\mathbb{Z}$,
$\nu^{(2l)}(W(n))<q_{1}^{-\rho_{2}c_{1}}.$ (5.16)
If we take
$L_{2}>\frac{12}{c_{1}},$ (5.17)
then (5.15) implies
$\nu^{(2l)}(A_{1}^{\prime\prime})>q_{1}^{-\frac{\rho_{2}c_{1}}{2}},$
which together with (5.16) implies the existence a set $A_{1}\subset A_{1}^{\prime\prime}$ satisfying all requirements in Lemma 5.7.
5.2. A large set of commutative matrices
Our goal in this section is to prove
Lemma 5.18.
Let $A_{1}$ be as given in Lemma 5.7. There is an element $\alpha_{0}\in A_{1}$ and a set $A_{2}\subset\mathbb{P}_{1}(A)^{2}$ such that
  1. (1)
    $\forall\gamma_{1},\gamma_{2}\in A_{2}$, $\gamma_{1}\neq\gamma_{2}\Rightarrow\gamma_{1}\neq\gamma_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}\}})$,
  2. (2)
    $|A_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{3}},$ (5.19)
  3. (3)
    $A_{2}$ commutes with $\alpha_{0}$ mod $q_{1}^{\{\rho_{2}\}}$.
Proof.
Denote
$Q_{0}=q_{1}^{\{\rho_{2}\}}.$
First, it follows from Proposition 4.1 that for any $\gamma_{0}\in\Gamma_{q}$,
$\nu^{(2l)}\{\gamma\in\Gamma:\gamma\equiv\gamma_{0}(\operatorname{mod}Q_{0})\}<Q_{0}^{-c_{1}}.$ (5.20)
Therefore,
$|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|>Q_{0}^{c_{1}}q^{-3\delta}>Q_{0}^{\frac{c_{1}}{2}},$ (5.21)
recalling (5.17).
Next, we use our assumption (5.2) to get a control over the growth of $\pi_{Q_{0}}(\mathbb{P}_{1}(A))$. By Lemma 2.2 from [Hel08], (5.2) implies for any $l\geq 3$,
$|\pi_{q}(A^{l})|\leq\left(\frac{|\pi_{q}(A\cdot A\cdot A)|}{|\pi_{q}(A)|}\right)^{l-2}|\pi_{q}(A)|$ (5.22)
$<$ $|\pi_{q}(A)|^{1+\delta(l-2)}.$ (5.23)
This implies
$|\pi_{Q_{0}}(\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A))|<|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{36}}.$ (5.24)
To see this, suppose not, then
$|\pi_{q}(A^{4})|\geq$ $|\pi_{Q_{0}}(\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A))|\cdot\max_{\gamma_{0}\in\Gamma_{Q_{0}}}|(\gamma_{1},\gamma_{2})\in A:\gamma_{1}\equiv\gamma_{0}(\operatorname{mod}Q_{0})|$
$\geq$ $|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{36}}\cdot\max_{\gamma_{0}\in\Gamma_{Q_{0}}}|(\gamma_{1},\gamma_{2})\in A:\gamma_{1}\equiv\gamma_{0}(\operatorname{mod}Q_{0})|$
$\geq$ $|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{\frac{c_{1}}{36}}\cdot|\pi_{q}(A)|$